# Do Green Reserve Assets Impact Stablecoin Stability?

### 25 Mar 2026 v1

@@ 4 Conclusion and Policy Implications
The empirical investigation presented in this study provides a comprehensive taxonomy of the “Climate-Liquidity Nexus,” offering the first longitudinal analysis of how the transition to green-collateralized reserves has structurally altered the stability of the digital dollar. Through the application of high-frequency VECM and GARCH(1,1) frameworks, we have demonstrated that the mandate established by the 2025 GENIUS Act has inadvertently traded the “information-insensitivity” of the stablecoin peg for the systemic volatility inherent in the nascent green bond market. My results indicate a profound regime shift; the digital monetary system has transitioned from a framework of exogenous stability (backed by the deep liquidity of US Treasuries) to one of endogenous fragility, where the unit of account is now a synthetic derivative of climate-risk-weighted assets.
The most critical finding of this research is the quantification of the “Liquidity Hysteresis” gap. We have proven that the recovery half-life of a green-backed peg is 5.4 times longer than that of its traditional predecessors. This extension of the de-pegging duration from hours to over 15 days represents a catastrophic failure for an asset intended to serve as a high-velocity settlement instrument. In the interconnected architecture of Decentralized Finance (DeFi), a fifteen-day de-pegging event is not a localized failure; it is a catalyst for cross-protocol liquidations, oracle failures, and a total evaporation of confidence in the “dollar-equivalent” status of the stablecoin. The 45% variance capture by green bonds identified in my FEVD analysis suggests that the stablecoin issuer has effectively ceded control over their monetary policy to the idiosyncratic shocks of the green secondary market.
@@
In light of these findings, we propose a rigorous three-pillar amendment to the GENIUS Act and broader stablecoin regulation. First, regulators must implement a “Liquidity Match Mandate,” requiring that any reserve asset with a secondary market turnover below a specific threshold (such as green municipal bonds) must be 1:1 matched with overnight reverse-repo facilities. Second, the current Value-at-Risk (VaR) standards for reserve adequacy must be abandoned in favor of Extreme Value Theory (EVT) models that explicitly account for the “fat tails” and negative skew documented in my density estimations. Finally, we call for the establishment of a specialized “Green Liquidity Facility” (GLF) at the central bank level. This facility would act as a buyer of last resort for green bonds during physical risk events, preventing the fire-sale feedback loops that characterized the 2025 crisis.
@@ 5 VECM and GARCH Specification
The primary econometric engine of this study is the Vector Error Correction Model (VECM), which allows for the simultaneous estimation of long-run equilibrium and short-run dynamic adjustment. We begin with a $p$-order Vector Autoregressive (VAR) model in levels:
@@
where $y_{t}=[Peg_{t},Green_{t}]^{\prime}$ is the vector of endogenous variables, representing the stablecoin peg and the green bond index respectively. By subtracting $y_{t-1}$ from both sides and rearranging terms, we derive the structural VECM representation:
@@
To address the observed heteroskedasticity and volatility clustering in the high-frequency residuals, we specify the conditional variance $\sigma^{2}_{t}$ of $\epsilon_{t}$ using a GARCH(1,1) process:
@@
where $\omega>0$ and $a,b\geq 0$. The persistence of volatility shocks is measured by $a+b$. In the post-GENIUS Act regime, we identify that $a+b\to 1$, indicating that climate-induced volatility is not transitory but lingers within the digital asset ecosystem.

### 28 Mar 2026 v2

@@ 4 Conclusion and Policy Implications
The empirical investigation presented in this study provides a comprehensive taxonomy of the “Climate-Liquidity Nexus,” offering the first longitudinal analysis of how the transition to green-collateralized reserves has structurally altered the stability of the digital dollar. Through the application of high-frequency VECM and GARCH(1,1) frameworks, I have demonstrated that the mandate established by the 2025 GENIUS Act has inadvertently traded the “information-insensitivity” of the stablecoin peg for the systemic volatility inherent in the nascent green bond market. My results indicate a profound regime shift; the digital monetary system has transitioned from a framework of exogenous stability (backed by the deep liquidity of US Treasuries) to one of endogenous fragility, where the unit of account is now a synthetic derivative of climate-risk-weighted assets.
The most critical finding of this research is the quantification of the “Liquidity Hysteresis” gap. I have proven that the recovery half-life of a green-backed peg is 5.4 times longer than that of its traditional predecessors. This extension of the de-pegging duration from hours to over 15 days represents a catastrophic failure for an asset intended to serve as a high-velocity settlement instrument. In the interconnected architecture of Decentralized Finance (DeFi), a fifteen-day de-pegging event is not a localized failure; it is a catalyst for cross-protocol liquidations, oracle failures, and a total evaporation of confidence in the “dollar-equivalent” status of the stablecoin. The 45% variance capture by green bonds identified in my FEVD analysis suggests that the stablecoin issuer has effectively ceded control over their monetary policy to the idiosyncratic shocks of the green secondary market.
@@
In light of these findings, I propose a rigorous three-pillar amendment to the GENIUS Act and broader stablecoin regulation. First, regulators must implement a “Liquidity Match Mandate,” requiring that any reserve asset with a secondary market turnover below a specific threshold (such as green municipal bonds) must be 1:1 matched with overnight reverse-repo facilities. Second, the current Value-at-Risk (VaR) standards for reserve adequacy must be abandoned in favor of Extreme Value Theory (EVT) models that explicitly account for the “fat tails” and negative skew documented in my density estimations. Finally, I call for the establishment of a specialized “Green Liquidity Facility” (GLF) at the central bank level. This facility would act as a buyer of last resort for green bonds during physical risk events, preventing the fire-sale feedback loops that characterized the 2025 crisis.
@@ 5 VECM and GARCH Specification
The primary econometric engine of this study is the Vector Error Correction Model (VECM), which allows for the simultaneous estimation of long-run equilibrium and short-run dynamic adjustment. I begin with a $p$-order Vector Autoregressive (VAR) model in levels:
@@
where $y_{t}=[Peg_{t},Green_{t}]^{\prime}$ is the vector of endogenous variables, representing the stablecoin peg and the green bond index respectively. By subtracting $y_{t-1}$ from both sides and rearranging terms, I derive the structural VECM representation:
@@
To address the observed heteroskedasticity and volatility clustering in the high-frequency residuals, I specify the conditional variance $\sigma^{2}_{t}$ of $\epsilon_{t}$ using a GARCH(1,1) process:
@@
where $\omega>0$ and $a,b\geq 0$. The persistence of volatility shocks is measured by $a+b$. In the post-GENIUS Act regime, I identify that $a+b\to 1$, indicating that climate-induced volatility is not transitory but lingers within the digital asset ecosystem.